Selenium-83 has a half life of 25. 0 minutes. How many minutes would it take for a 10.0 mg sample to decay and have only 1.25 mg. remain?


n = ?

T = ? minutes


1
Expert's answer
2021-04-20T16:49:22-0400

"n=n_0(\frac{1}{2})^{\frac{T}{T_{1/2}}}"


"1.25mg=10.0mg\times(\frac{1}{2})^{\frac{T}{25.0min}}"


Solving for T,

T = 75.0 minutes


Answer: 75.0 minutes